How many grams of H2O will be formed when 26.5 g H2 reacts with 26.5 g of O2 according to the reaction:
2H2 + O2 → 2H2O
1
Expert's answer
2020-03-23T10:20:55-0400

We have the equation

2H2 + O2 → 2H2O.

Calculate how many moles of H2 and O2 there are in 26.5 g of each substance:


"n_{H_2}=\frac{m}{\Mu_{H_2}}=\frac{26.5}{2}=13.3\text{ mol},\\
\space\\
n_{O_2}=\frac{m}{\Mu_{O_2}}=\frac{26.5}{32}=0.828\text{ mol}."


We will calculate the reaction by oxygen because it has the smallest amount of substance. The amount of substance of water required is


"n_{H_2O}=2n_{O_2},\\
\space
\\m_{H_2O}=n_{H_2O}\cdot\Mu_{H_2O}=\\=2n_{O_2}\cdot\Mu_{H_2O}=\\
=2\cdot0.828\cdot18=29.8\text{ g}."